A Level Further Mathematics 9231 Paper 41, May/June 2024

Question paper and mark scheme50 marks7 questions

Solve it on ExamPal

Practise

Untimed. Write on the paper and check as you go.

Start

Exam

Timed, with a clock you set, like sitting the real paper.

Start

Opens full screen.

Papers open in

Opening the question paper

Read as textThe questions and the mark scheme as plain text. Easier to read on a phone.

Question paper

9231/41 · May/June 2024

Download PDF
1

The times taken by members of a large cycling club to complete a cross-country circuit have a normal distribution with mean μ minutes. The times taken, x minutes, are recorded for a random sample of 14 members of the club. The results are summarised as follows, where xˉ is the sample mean. xˉ=42.8Σ(x−xˉ)2=941.5 Find a 95% confidence interval for μ.

[4]
2

A large number of students are taking a Physics course. They are assessed by a practical examination and a written examination. The marks out of 100 obtained by a random sample of 15 students in each of the examinations are as follows.

StudentABCDEFGHIJKLMNO
Practical examination666324525976885148369172686760
Written examination635739504771876556397870616270

Use a sign test, at the 10% significance level, to test whether, on average, the practical examination marks are higher than the written examination marks.

[5]
3

A factory produces metal discs. The manager claims that the diameters of these discs have a median of 22.0 mm. The diameters, in mm, of a random sample of 12 discs produced by this factory are as follows.

22.420.922.821.523.222.923.921.719.823.622.623.0
a

Carry out a Wilcoxon signed-rank test, at the 10% significance level, to test whether there is any evidence against the manager's claim.

[7]
b

State an assumption that is necessary for this test to be valid.

[1]
4

The random variable Y is the sum of two independent observations of the random variable X. The probability generating function GY(t) of Y is given by GY(t)=t2(4−3t)4.

a

Find E(Y).

[3]
b

Write down an expression for the probability generating function of X.

[1]
c

Find P(X=4).

[3]
5

Two companies, P and Q, produce a certain type of paint brush. An independent examiner rates the quality of the brushes produced as poor, satisfactory or good. He takes a random sample of brushes from each company. The examiner's ratings are summarised in the table.

CompanyPoorSatisfactoryGood
P184364
Q222231
a

Test, at the 5% significance level, whether quality of brushes is independent of company.

[7]
b

Compare the quality of the brushes produced by the two companies.

[1]
6

Jade is a swimming instructor at a sports college. She claims that, as a result of an intensive training course, the mean time taken by students to swim 50 metres has reduced by more than 1 second. She chooses a random sample of 10 students. The times taken, in seconds, before and after the training course are recorded in the table.

StudentABCDEFGHIJ
Time before course54.247.452.159.055.351.048.952.258.451.4
Time after course50.146.352.558.851.448.449.548.758.351.4
a

Test, at the 10% significance level, whether Jade's claim is justified.

[7]
b

State an assumption that is necessary for this test to be valid.

[1]
7

The continuous random variable X has probability density function f given by f(x)={x4(4−x2)0⩽x⩽2,0otherwise.

a

Find Var(X).

[4]
b

The continuous random variable Y is defined by Y=X2. Find the probability density function of Y.

[4]
c

Find the exact value of the median of Y.

[2]

Mark scheme

9231/41 · May/June 2024

Download PDF

1 · 4 marks

  • B1

    s2=941.513

  • M1

    CI: 42.8±2.160s214

  • A1

    2.160 seen in a CI formula

  • A1

    [37.9,47.7]

2 · 5 marks

  • B1

    H0: practical results and written results are equal H1: practical results are greater than written results

  • M1

    There are 10 + signs; B(15,0.5) soi

  • A1

    (P(X⩾10)=1−P(X⩽9)=)1−0.849=0.151

  • M1

    Compare with 0.1: '0.151' > 0.1 accept H0

  • A1

    Insufficient evidence to suggest that practical results are greater than written results

3a · 7 marks

  • M1

    Signed differences 0.4 -1.1 0.8 -0.5 1.2 0.9 1.9 -0.3 -2.2 1.6 0.6 1.0

  • M1

    Ranks: 2 -8 5 -3 9 6 11 -1 -12 10 4 7

  • A1

    (P+=54) P−=24

  • B1

    H0: population median = 22.0 H1: population median ≠ 22.0

  • B1

    Critical value from table = 17

  • M1

    '24' > '17' Accept H0

  • A1

    Insufficient evidence to support (population) median not being 22.0 Insufficient evidence against the manager’s claim

3b · 1 mark

  • B1

    Distribution is symmetric about the population median Underlying/population distribution is symmetric about the median

4a · 3 marks

  • M1

    GY(t)=t2(4−3t)4, GY′(t)=(4−3t)4(2t)−−12t2(4−3t)3(4−3t)8 or 2t(4−3t)4+12t2(4−3t)5

  • M1

    E(Y)=GY′(1)=14

  • A1

    14

4b · 1 mark

  • B1

    GX(t)=t(4−3t)2

4c · 3 marks

  • M1

    GX(t)=t(4−3t)−2=t16(1−34t)−2=t16(1+32t−3×916t2+4×2764t3+… )

  • M1

    Use P(X=4)=their coefficient of t4

  • A1

    27256

5a · 7 marks

  • M1

    Calculate E values

  • A1

    Expected values table

  • M1

    Chi-squared contributions

  • A1

    Test statistic = 6.56

  • B1

    H0: Quality (of items) is independent of company H1: Quality (of items) is not independent of company

  • M1

    6.56 > 5.991 Reject H0

  • A1

    There is sufficient evidence to suggest that the quality of brushes is not independent of company.

5b · 1 mark

  • B1

    P produces better quality brushes (than Q) or Q produces worse quality brushes (than P) P has fewer poor quality brushes than expected AND Q has more poor quality brushes than expected

6a · 7 marks

  • M1

    Signed differences: 4.1 1.1 -0.4 0.2 3.9 2.6 -0.6 3.5 0.1 0

  • M1

    ∑d=14.5, ∑d2=52.81 s2=19(52.81−14.5210)=3.532

  • B1

    H0:μB−μA=1 H1:μB−μA>1

  • M1

    t=1.45−1s210

  • A1

    0.757

  • M1

    Critical value is 1.383: ‘0.757’ < 1.383 Accept H0

  • A1

    Insufficient evidence to support Jade’s claim

6b · 1 mark

  • B1

    Distribution of population differences is normal

7a · 4 marks

  • B1

    f(x)={x4(4−x2)0⩽x⩽2,0otherwise. (E(X)=)∫0214(4x2−x4) dx=14[43x3−15x5]02=1615 (=1.067)

  • M1

    (E(X)=)∫0214(4x1.5−x3.5) dx=14[85x2.5−29x4.5]02[=32452=1.005(7)]

  • M1

    Var(X)=E(X)−(E(X))2 used

  • A1

    [1615−(32452)2]=0.0553

7b · 4 marks

  • B1

    F(x)=116(8x2−x4) [0⩽x⩽2]

  • M1

    G(y)=116(8y−y2) [0⩽y⩽4]

  • M1

    g(y)=116(8−2y) [0⩽y⩽4]

  • A1

    g(y)={116(8−2y)0⩽y⩽4,0otherwise.

7c · 2 marks

  • M1

    116(8m−m2)=12

  • A1

    [m2−8m+8=0,]m=4−22

Questions

  1. 1
    Question 14 marks
    Mark scheme ›
  2. 2
    Question 25 marks
    Mark scheme ›
  3. 3
    Question 38 marks
    Mark scheme ›
  4. 4
    Question 47 marks
    Mark scheme ›
  5. 5
    Question 58 marks
    Mark scheme ›
  6. 6
    Question 68 marks
    Mark scheme ›
  7. 7
    Question 710 marks
    Mark scheme ›

More 9231 papers

Other papers from May/June 2024

Solve it on ExamPalStart